# T2ICount: Enhancing Cross-modal Understanding for Zero-Shot Counting

### v1

@@ Abstract
Zero-shot object counting aims to count instances of arbitrary object categories specified by text descriptions. Existing methods typically rely on vision-language models like CLIP, but often exhibit limited sensitivity to text prompts. We present T2ICount, a diffusion-based framework that leverages rich prior knowledge and fine-grained visual understanding from pretrained diffusion models. While one-step denoising ensures efficiency, it leads to weakened text sensitivity. To address this challenge, we propose a Hierarchical Semantic Correction Module that progressively refines text-image feature alignment, and a Representational Regional Coherence Loss that provides reliable supervision signals by leveraging the cross-attention maps extracted from the denosing U-Net. Furthermore, we observe that current benchmarks mainly focus on majority objects in images, potentially masking models’ text sensitivity. To address this, we contribute a challenging re-annotated subset of FSC147 for better evaluation of text-guided counting ability. Extensive experiments demonstrate that our method achieves superior performance across different benchmarks. Code is available at https://github.com/cha15yq/T2ICount.
@@ 1 Introduction
In this work, we propose a framework that leverages single-step features from diffusion models to achieve zero-shot counting. Though computationally efficient, this design sacrifices the text awareness that diffusion models typically build through multiple denoising steps [31], resulting in limited text sensitivity. To overcome this limitation, we propose a new Hierarchical Semantic Correction Module (HSCM) to compensate for the weakened text-image interaction. The HSCM progressively rectifies the semantic-visual discrepancy through multi-scale feature rectification. We complement this with a novel Representational Regional Coherence Loss ($\mathcal{L}_{\text{RRC}}$) that enhances cross-modal alignment. $\mathcal{L}_{\text{RRC}}$ leverages cross attention maps from the diffusion model to delineate the general foreground regions, thereby solving a key challenge where only point-level annotations are available for supervision: while positive samples can be determined through density thresholds, identifying reliable negative regions is difficult without instance-level annotations. By capturing general object shapes, $\mathcal{L}_{\text{RRC}}$ enables better positive-negative sample selection for more precise feature learning.
@@ 2 Related Work
Figure 2: Overview of the proposed T2ICount. Our method is based on single denoising step. An input image and text prompts specifying the category to be counted are fed into the pre-trained text-to-image diffusion model. Feature maps extracted from the decoder of the U-Net are passed through the Hierarchical Semantic Correction Module to enhance textual awareness, producing the final features used to estimate the density map. Text-image similarity maps are generated at intermediate stages and are supervised by the Representational Regional Coherence Loss. The ground-truth density map and the fused cross attention maps ($\widehat{\mathcal{A}}^{cross}$) are used to generate the positive-negative-ambiguous (PNA) map, providing supervision signals for this loss.
@@ 2.1 Few-shot Object Counting
2.2 Reference-less Object Counting
Reference-less object counting has been proposed as an alternative which counts instances of the the dominant class in an image without any exemplars. In pioneering work, RepRPN-Counter [22] proposes a module that selects the most frequently repeating objects from the image as a substitute for exemplars. RCC [4] offers a solution using a pre-trained ViT to directly regress object counts with weak image-level supervision. Despite this progress, a key limitation of reference-less counting is the inability to specify which objects to count, particularly in complex scenes containing multiple object types.
@@ 2.3 Zero-shot Object Counting
Zero-shot object counting addresses this issue by incorporating external information, such as text descriptions, allowing the model to count specific object categories without exemplars. Xu et al. [32] propose the first method for this task, building on a few-shot counting framework in which a text-conditioned variational autoencoder (VAE) is used to generate visual exemplars specified by the text. Later, three concurrent workshave leveraged the association between text and image embeddings learned by the pre-trained language-vision foundation model, CLIP [20], for text-guided zero-shot counting. These are CLIP-Count [6], VLCounter [7], and CountX [1], respectively. All three methods demonstrate similar counting performance but remain far behind state-of-the-art visual exemplar-based frameworks. More recently, new methods have been proposed that attempt to adapt different vision-language models for this task. PseCo [36] introduces a framework that leverages the Segment Anything model [8] for proposal generation and uses CLIP for classification based on text specification. VA-Count [37] leverages Grounding DINO [12] for initial text-specified detection, then proposes a mechanism to select good visual exemplars from its predictions for match-based counting. Despite rapid progress, a key challenge remains in assessing how well the model is counting the objects specified by the text, largely due to the characteristics of the FSC-147 dataset, in which all images are annotated with only a single object class, and in most cases this is the majority class. To overcome this, we propose a simple synthetic dataset specifically designed to evaluate the model’s behavior in text-guided counting.
@@ 3.1 Text Insensitivity in Single-Step Denoising
Diffusion models have emerged as a powerful family of generative models that learn complex data distributions through a gradual denoising process. In our framework, we leverage Stable Diffusion [24], which performs diffusion in a compact latent space in which images and text are encoded via a VAE encoder and CLIP encoder respectively. Formally, given a latent variable $z_{t}$ corrupted from the compressed image representation $z_{0}$ through a forward diffusion process:
@@ 3.3 Representational Regional Coherence Loss
$p_{jk}=\left\{\begin{aligned} 1,&\quad\text{if}\quad D^{\text{gt}}_{jk}\geq\tau,\\ 0,&\quad\text{else if}\quad{\mathcal{A}}^{\text{cross}}_{jk}\leq 0.5\times\widehat{\mathcal{A}}^{\text{cross}}_{jk},\\ -1,&\quad\text{otherwise}.\end{aligned}\right.$ (6)
Here, $\tau$ is a threshold, $\widehat{\mathcal{A}}^{\text{cross}}$ represents the average value of ${\mathcal{A}}^{\text{cross}}$. We visualize some intermediate results in our supervision signal generation in Fig 3: (d) ${\mathcal{A}}^{\text{cross}}$, (e) pseudo-background maps obtained by directly binarizing the ${\mathcal{A}}^{\text{cross}}$ using the second condition in Eq 6, and (f) PNA maps. Note that these ambiguous regions (shown in gray) are commonly treated as negative regions in traditional methods.

### v2

@@ Abstract
Zero-shot object counting aims to count instances of arbitrary object categories specified by text descriptions. Existing methods typically rely on vision-language models like CLIP, but often exhibit limited sensitivity to text prompts. We present T2ICount, a diffusion-based framework that leverages rich prior knowledge and fine-grained visual understanding from pretrained diffusion models. While one-step denoising ensures efficiency, it leads to weakened text sensitivity. To address this challenge, we propose a Hierarchical Semantic Correction Module that progressively refines text-image feature alignment, and a Representational Regional Coherence Loss that provides reliable supervision signals by leveraging the cross-attention maps extracted from the denoising U-Net. Furthermore, we observe that current benchmarks mainly focus on majority objects in images, potentially masking models’ text sensitivity. To address this, we contribute a challenging re-annotated subset of FSC147 for better evaluation of text-guided counting ability. Extensive experiments demonstrate that our method achieves superior performance across different benchmarks. Code is available at https://github.com/cha15yq/T2ICount.
@@ 1 Introduction
In this work, we propose a framework that leverages single-step features from diffusion models to achieve zero-shot counting. Though computationally efficient, this design sacrifices the text awareness that diffusion models typically build through multiple denoising steps [32], resulting in limited text sensitivity. To overcome this limitation, we propose a Hierarchical Semantic Correction Module (HSCM) to compensate for the weakened text-image interaction. The HSCM progressively rectifies the semantic-visual discrepancy through multi-scale feature rectification. We complement this with a novel Representational Regional Coherence Loss ($\mathcal{L}_{\text{RRC}}$) that enhances cross-modal alignment. $\mathcal{L}_{\text{RRC}}$ leverages cross attention maps from the diffusion model to delineate the general foreground regions, thereby solving a key challenge where only point-level annotations are available for supervision: while positive samples can be determined through density thresholds, identifying reliable negative regions is difficult without instance-level annotations. By capturing general object shapes, $\mathcal{L}_{\text{RRC}}$ enables better positive-negative sample selection for more precise feature learning.
@@ 2 Related Work
Figure 2: Overview of the proposed T2ICount. Our method is based on single denoising step. An input image and text prompts specifying the category to be counted are fed into the pre-trained text-to-image diffusion model. Feature maps extracted from the decoder of the U-Net are passed through the Hierarchical Semantic Correction Module to enhance textual awareness, producing the final features used to estimate the density map. Text-image similarity maps are generated at intermediate stages and are supervised by the Representational Regional Coherence Loss. The ground-truth density map and the fused cross attention maps ($\widehat{\mathcal{A}}^{cross}$) are used to generate the positive-negative-ambiguous (PNA) map, providing supervision signals for this loss. In the training process, the VAE encoder and the text encoder are frozen while the U-Net and HSCM are being trained.
@@ 2.2 Zero-shot Object Counting
Zero-shot object counting addresses this issue by incorporating external information, such as text descriptions, allowing the model to count specific object categories without exemplars. Xu et al. [33] propose the first method for this task, building on a few-shot counting framework in which a text-conditioned variational autoencoder (VAE) is used to generate visual exemplars specified by the text. Later, three concurrent workshave leveraged the association between text and image embeddings learned by the pre-trained language-vision foundation model, CLIP [21], for text-guided zero-shot counting. These are CLIP-Count [7], VLCounter [8], and CountX [1], respectively. All three methods demonstrate similar counting performance but remain far behind state-of-the-art visual exemplar-based frameworks. More recently, new methods have been proposed that attempt to adapt different vision-language models for this task. PseCo [38] introduces a framework that leverages the Segment Anything model [9] for proposal generation and uses CLIP for classification based on text specification. VA-Count [39] leverages Grounding DINO [13] for initial text-specified detection, then proposes a mechanism to select good visual exemplars from its predictions for match-based counting. Despite rapid progress, a key challenge remains in assessing how well the model is counting the objects specified by the text, largely due to the characteristics of the FSC-147 dataset, in which all images are annotated with only a single object class, and in most cases this is the majority class. To overcome this, we re-annotated a portion of the FSC-147 dataset specifically designed to evaluate the model’s behavior in text-guided counting.
@@ 3.1 Text Insensitivity in Single-Step Denoising
Diffusion models have emerged as a powerful family of generative models [37, 3] that learn complex data distributions through a gradual denoising process. In our framework, we leverage Stable Diffusion [25], which performs diffusion in a compact latent space in which images and text are encoded via a VAE encoder and CLIP encoder respectively. Formally, given a latent variable $z_{t}$ corrupted from the compressed image representation $z_{0}$ through a forward diffusion process:
@@ 3.3 Representational Regional Coherence Loss
$p_{jk}=\left\{\begin{aligned} 1,&\quad\text{if}\quad D^{\text{gt}}_{jk}\geq\tau,\\ 0,&\quad\text{else if}\quad{\mathcal{A}}^{\text{cross}}_{jk}\leq\theta,\\ -1,&\quad\text{otherwise}.\end{aligned}\right.$ (6)
Here, $\tau$ and $\theta$ are thresholds. We visualize some intermediate results in our supervision signal generation in Fig 3: (d) ${\mathcal{A}}^{\text{cross}}$, (e) pseudo-background maps obtained by directly binarizing the ${\mathcal{A}}^{\text{cross}}$ using the second condition in Eq 6, and (f) PNA maps. Note that these ambiguous regions (shown in gray) are commonly treated as negative regions in traditional methods.
@@ 5 Conclusion
Acknowledgements
This work was supported by the Biotechnology and Biological Sciences Research Council (grant number BB/Y513908/1). This work was also funded by the National Natural Science Foundation of China (62376070, 62076195).